# Right for Right Reasons: Large Language Models for Verifiable Commonsense Knowledge Graph Question Answering

### v1

@@ 6 Limitations
In this section, we honestly and faithfully present the limitations of our current research and briefly discuss the possible future directions. The quality of the reasoning process in our methodology relies on the quality of the axioms generated. We observe through our experiments that in cases where the quality of axioms is insufficient, the reasoner is misled resulting in an undetermined answer at the end of the exploration budget identified. It is necessary to mention that in our experiments, we used $b=2$, a relatively small value for the maximum breadth since we expected the model to obtain an answer in a rather small number of efforts. Furthermore, due to the importance of avoiding hallucination, our model takes a conservative and rigorous approach to ground every factual premise on KG triples. Therefore, our model typically leaves more questions unanswered than other baselines. Obviously, we considered the unanswered questions as an incorrect response in calculating the accuracy.

### v2

@@ 4.4.2 Claim Verification
A statistical analysis of these results is provided in Appendix B, which verifies that $R^{3}$ statistically significantly reduces sources of hallucination on three of the studied datasets. We also provide anecdotal examples of $R^{3}$’s performance in addressing LLM misbeliefs in Appendix C.
@@ 6 Limitations
While we believe this work has made significant forward progress in leveraging KG content for commonsense question answering (QA), our method $R^{3}$ (like any QA method) has natural limitations that we hope will encourage further investigation and future work.
The quality of the reasoning process in $R^{3}$ relies on the quality of the natural language axioms generated. We observe through our experiments that in cases where the quality of axioms is insufficient, the reasoner is misled resulting in an undetermined answer at the end of the exploration budget identified. Furthermore, due to the importance of avoiding hallucination, our model takes a conservative and rigorous approach to ground every factual premise on KG triples. Therefore, our model typically leaves more questions unanswered than other baselines (which we considered an incorrect response in calculating the accuracy).